### v1

@@ 4.3.2 Pre-training of SpecFormer
To enable more effective encoding of molecular spectra, we introduce a masked reconstruction pre-training objective. In addition, to overcome the scarcity of spectral pre-training data by leveraging large-scale molecular structure pre-training, we incorporate a contrastive learning objective to align spectral and structural representations. The complete pre-training framework is illustrated in Figure 1(C).
Masked patches reconstruction pre-training for spectra. To ensure that the spectrum encoder can effectively extract and represent information from molecular spectra, we adopt a masked patches reconstruction (MPR) pre-training strategy. Inspired by the effectiveness of masked reconstruction across multiple domains [54, 55, 56, 57, 58, 59], MPR guides the learning of SpecFormer by encouraging it to reconstruct masked portions of spectral data.
After segmenting the spectra into patches, we randomly mask a fraction of these patches—according to a predefined ratio $\alpha$—by replacing them with zero vectors. The masked patch sequences then undergo patch and position encoding, which conceals their semantic content (such as absorption intensities at certain wavelengths) but retains their positional information, aiding the reconstruction task.
Once processed by SpecFormer, the encoded representations corresponding to the masked patches are passed through a reconstruction head specific to each spectrum. The original values of the masked patches are then predicted, with the mean squared error (MSE) between the reconstructed and true patch values serving as the training objective:
@@
Contrastive learning between 3D structures and spectra. To align spectral and 3D molecular representations, we introduce a contrastive learning objective in addition to the MPR. The 3D embeddings are learned under the guidance of a denoising objective [60, 31]. Here, the 3D embedding ${\bm{z}}_{x}\in\mathbb{R}^{d}$ and spectra embedding ${\bm{z}}_{s}\in\mathbb{R}^{d}$ of the same molecule are treated as positive pairs, while all other pairings are considered negative. The contrastive objective is designed to maximize similarity between positive pairs while minimizing similarity with negatives, using the InfoNCE loss [61]:
@@
Two-stage pre-training pipeline. Although spectral datasets are scarce, there exists a wealth of large-scale unlabeled molecular structure datasets. By leveraging our proposed contrastive alignment framework between spectra and structures, we can transfer knowledge from large-scale pre-training on molecular structures to enhance the learning of SpecFormer. To fully exploit both spectral and structural information, we adopt a two-stage training protocol: the first stage performs pre-training on a large dataset [62] containing only 3D structures using the denoising objective, while the second stage utilizes a dataset with available spectra to jointly optimize the overall objective:
@@ Appendix B Datasets
We conduct experiments on the QM9S dataset [8, 31], which augments the original QM9 [63] molecular dataset with simulated spectra. For each molecule, we extract its structure and corresponding IR, Raman, and UV-Vis spectra, forming a multi-modal spectral condition for our structure elucidation task. The dataset contains 133,885 molecules in total. Following the original setting in DetaNet [8], we adopt the same training, validation, and test split strategy, which partitions the dataset into 90% for training, 5% for validation, and 5% for testing.
@@ Appendix C Metrics
Aiming to generating chemically valid and complete molecules, modeling accurate molecular distributions, and achieving achieve accurate sturcture elucidation from spectra, we elaborately design a series of evaluation metrics to reflect the generation quality.
@@ C.1 Basic Metrics for Molecular Generation
The primary consideration is the chemical rationality of the generated molecular structures. The Validity metric ensures that the molecular structures conform to basic chemical rules, including chemically plausible valences (e.g., carbon atoms typically form four covalent bonds), and the overall bonding patterns are chemically reasonable (e.g., avoiding unrealistic bonds between atoms, maintaining aromatic delocalization). Building upon this, we introduce the Validity and Complete (V&C) indicator, which not only requires chemical validity but also guarantees that the molecular graph forms a single connected entity instead of fragmented parts. This is particularly important for evaluating the model’s ability to generate fully connected, functional molecules.
To further assess the model’s capacity for innovation, we measure the proportions of molecules satisfying Validity and Unique (V&U) and Validity, Unique, and Novelty (V&U&N) criteria. Uniqueness is determined by canonicalizing molecular representations and removing duplicates, thus ensuring diversity among the generated structures. Novelty is evaluated by comparing the generated molecules against the training set, reflecting the model’s ability to produce genuinely novel chemotypes rather than simply memorizing training data.
@@
To evaluate the model’s ability to learn the data distribution of real molecules, we adopt several distribution-based metrics. The Fréchet ChemNet Distance (FCD) measures the similarity between the distributions of generated molecules and reference molecules in a high-dimensional learned feature space. The Similarity to Nearest Neighbor (SNN) metric assesses the representativeness of the generated set by measuring the Tanimoto similarity between each generated molecule and its nearest neighbors in the test set, reflecting both diversity and coverage.
We additionally employ metrics based on molecular structural features: Fragment Similarity (Frag), based on BRICS decomposition, evaluates distributional alignment at the functional group level; while Scaffold Similarity (Scaf), using Bemis–Murcko scaffold analysis, quantifies similarity and diversity at the core-structure level. These complementary metrics jointly reflect the model’s ability to capture both local functional groups and global structural patterns, thereby characterizing the complexity of the generated molecular structures.
3D Geometry-based Structural Evaluation
Beyond chemical rules in 2D representations, assessing the 3D geometric consistency of generated molecules is critical, especially for downstream applications such as molecular docking or molecular structure elucidation from spectra. We therefore introduce a set of geometric evaluation metrics based on fundamental molecular geometry features: bond lengths, bond angles, and dihedral angles.
Bond evaluation compares the distributions of bond lengths in the generated set against those in the reference set (test set molecules), capturing whether typical interatomic distances are preserved. Angle evaluation similarly assesses the distributions of bond angles, reflecting the local spatial arrangements of connected atoms and ring systems. Dihedral evaluation analyzes the distributions of torsional (dihedral) angles, providing insights into conformational consistency and stereochemical plausibility of the generated molecules.
To quantify the similarity between these geometric features of the generated and reference molecules, we employ maximum mean discrepancy (MMD), which measures the distance between two distributions in a kernel space. A lower MMD indicates that the generated molecules more faithfully reproduce realistic 3D geometries.
It is important to note that for all distribution-based and geometric metrics, the test set serves as the reference, while the novelty metric uses the training set as the reference to measure whether generated molecules are unseen. This distinction ensures a clear separation between distributional fidelity and generative innovation.
@@ C.2 Metrics for Molecular Structure Elucidation
The Top-$K$ Accuracy quantifies the model’s ability to recover the exact target structure among its top-$K$ generated candidates. For each condition characterized by a set of spectra, the model is sampled $K$ times to generate $K$ molecular structure candidates. If any of the $K$ candidates exactly matches the ground-truth structure, the prediction is considered correct. Formally, the metric is defined as:
@@
Maximum Common Edge Subgraph (MCES)
To quantify graph-structural overlap, we adopt the Maximum Common Edge Subgraph (MCES) metric:
$\operatorname{MCES}=\mathbb{E}_{\mathcal{G}}\left[\max_{\mathcal{H}\subseteq\mathcal{G},\,\mathcal{H}\subseteq\hat{\mathcal{G}}}|E(\mathcal{H})|\right],$ (41)
where $E(\mathcal{H})$ denotes the set of edges in the common subgraph $\mathcal{H}$. This indicator captures the extent of exact subgraph matching between the target and predicted molecular graphs, with higher values reflecting a greater degree of structural recovery.
Fingerprint-based Similarity Metrics
@@
Fragment-based Similarity (Fraggle)
The Fraggle Similarity metric decomposes $\mathcal{G}$ into a set of chemically meaningful fragments =$\mathcal{F}=\{f_{1},f_{2},\dots,f_{n}\}$ by strategies including double acyclic or exocyclic bond cuts with specified fragment size constraints. For each fragment $f_{i}$, Fraggle computes two Tanimoto similarities over RDKit fingerprints between the target molecule $\mathcal{G}$ and the predicted molecule $\hat{\mathcal{G}}$: one based on the standard RDKit fingerprint over the entire molecule, and another based on a masked RDKit fingerprint where atoms outside $f_{i}$ with a Tversky similarity below 0.8 are replaced with wildcards. The fragment-level score is then taken as the maximum of these two values:
@@
Functional Group-based Similarity
Lastly, to evaluate functional group consistency, we define a set of chemically significant functional groups (e.g., alkane, alcohol, amine, carboxylic acid, etc.) described by SMARTS patterns. For a pair of molecules $\mathcal{G}$ (ground truth) and $\hat{\mathcal{G}}$ (prediction), we extract their respective functional group sets $FG(\mathcal{G})$ and $FG(\hat{\mathcal{G}})$. The Functional Group Similarity (FGSim) is then computed as:
$\operatorname{FGSim}=\mathbb{E}_{\mathcal{G}}\left[\frac{\left|FG(\mathcal{G})\cap FG(\hat{\mathcal{G}})\right|}{\left|FG(\mathcal{G})\cup FG(\hat{\mathcal{G}})\right|}\right].$ (47)

### v2

@@ 4.3.2 Pre-training of SpecFormer
To enable more effective encoding of molecular spectra, we introduce a masked-reconstruction pre-training objective. In addition, to mitigate the scarcity of spectral pre-training data by leveraging large-scale molecular structure pre-training, we incorporate a contrastive learning objective to align spectral and structural representations. The complete pre-training framework is illustrated in Figure 1(C).
Masked patches reconstruction pre-training for spectra. To ensure that the spectrum encoder can effectively extract and represent information from molecular spectra, we adopt a masked patches reconstruction (MPR) pre-training strategy. Inspired by the effectiveness of masked reconstruction across multiple domains [58, 59, 60, 61, 62, 63], MPR guides the learning of SpecFormer by encouraging it to reconstruct the masked portions of the spectral data.
After segmenting the spectra into patches, we randomly mask a fraction of these patches—according to a predefined ratio $\alpha$—by replacing them with zero vectors. The masked patch sequences then undergo patch and positional encoding, which conceals their semantic content (such as absorption intensities at certain wavelengths) but retains their positional information, thereby aiding the reconstruction task.
After processed by SpecFormer, the encoded representations corresponding to the masked patches are passed through a reconstruction head specific to each spectrum. The original values of the masked patches are subsequently predicted, with the mean squared error (MSE) between the reconstructed and true patch values serving as the training objective:
@@
Contrastive learning between 3D structures and spectra. To align spectral and 3D molecular representations, we introduce a contrastive learning objective in addition to the MPR. The 3D embeddings are learned under the supervision of a denoising objective [64, 30]. Here, the 3D embedding ${\bm{z}}_{x}\in\mathbb{R}^{d}$ and the spectra embedding ${\bm{z}}_{s}\in\mathbb{R}^{d}$ of the same molecule are treated as positive pairs, while all other pairings are considered negative. The contrastive objective is designed to maximize similarity between positive pairs while minimizing similarity with negatives, using the InfoNCE loss [65]:
@@
Two-stage pre-training pipeline. Although spectral datasets are scarce, a wealth of large-scale unlabeled molecular structure datasets exists. By leveraging our proposed contrastive alignment framework between spectra and structures, we transfer knowledge from large-scale pre-training on molecular structures to enhance the learning of SpecFormer. To fully exploit both spectral and structural information, we adopt a two-stage training protocol: the first stage performs pre-training on a large dataset [66] containing only 3D structures using the denoising objective, while the second stage utilizes a dataset with available spectra to jointly optimize the overall objective:
@@
5 Data availability
The preprocessed dataset can be accessed at https://huggingface.co/datasets/AzureLeon1/DiffSpectra.
6 Code availability
The source code is open-source and can be accessed at https://github.com/AzureLeon1/DiffSpectra. It is available for non-commercial use. Trained model checkpoints can be found at https://huggingface.co/AzureLeon1/DiffSpectra.
7 Acknowledgements
This work was supported in part by the National Science and Technology Major Project (2023ZD0120901), the Strategic Priority Research Program of Chinese Academy of Sciences (XDA0480102), the Ministry of Education (MOE T1251RES2309 and MOE T2EP20125-0039), the Agency for Science, Technology and Research (ASTAR H25J6a0034), and the Damo Academy through Damo Academy Research Intern Program.
8 Author Contributions Statement
L.W., Y.R., and T.X. conceived the initial idea for the project. L.W. and Z.Z. developed the model. L.W. processed the dataset and trained the model. L.W. and Z.Z. carried out the experiments using the trained model. L.W., T.X., and Z.Z. analyzed the results and created illustrations and data visualizations. L.W. drafted the initial manuscript. L.W., Y.R., Z.L, Q.L., Y.Z. participated in the revision of the manuscript. T.X. and Z.L. provided the instructions on AI modeling. Z.L., P.W., and Y.Z. provided the instructions on evaluation framework. The project was supervised by Y.R., D.Z., Q.L., S.W., L.W., Y.Z., with funding provided by Y.R., Q.L., S.W., and Y.Z.
9 Competing Interests Statement
The authors declare no competing interests.
@@ Appendix B Datasets
We conduct experiments on the QM9S dataset [8, 30], which augments the original QM9 [67] molecular dataset with simulated spectra. For each molecule, we extract its structure along with the corresponding IR, Raman, and UV-Vis spectra, forming a multi-modal spectral condition for our structure elucidation task. The dataset contains approximately 127k molecules in total. Following the dataset split protocol used in previous studies [38, 68, 45], we split QM9S into training, validation and test sets, which contain 97K, 17K and 13K samples respectively.
@@ Appendix C Metrics
Aiming to generating chemically valid and complete molecules, modeling accurate molecular distributions, and achieving achieve accurate structure elucidation from spectra, we elaborately design a series of evaluation metrics to reflect the generation quality.
@@ C.1 Basic Metrics for Molecular Generation
The primary consideration is the chemical rationality of the generated molecular structures. The Validity metric ensures that the molecular structures conform to basic chemical rules, including chemically plausible valences (e.g., carbon atoms typically form four covalent bonds), and the overall bonding patterns are chemically reasonable (e.g., avoiding unrealistic bonds between atoms, maintaining aromatic delocalization). To further assess the model’s capacity for innovation, we measure the proportions of molecules satisfying Validity and Unique (V&U) and Validity, Unique, and Novelty (V&U&N) criteria. Uniqueness is determined by canonicalizing molecular representations and removing duplicates, thus ensuring diversity among the generated structures. Novelty is evaluated by comparing the generated molecules against the training set, reflecting the model’s ability to produce genuinely novel molecules rather than simply memorizing training data.
@@
To evaluate the model’s ability to learn the data distribution of real molecules, we adopt several distribution-based metrics. The Fréchet ChemNet Distance (FCD) measures the similarity between the distributions of generated molecules and reference molecules in a high-dimensional learned feature space. The Similarity to Nearest Neighbor (SNN) metric assesses the representativeness of the generated set by measuring the Tanimoto similarity between each generated molecule and its nearest neighbors in the test set. We additionally employ metrics based on molecular structural features: Fragment Similarity (Frag), derived from BRICS decomposition, evaluates distributional alignment at the functional group level, whereas Scaffold Similarity (Scaf), based on Bemis–Murcko scaffold analysis, quantifies similarity and diversity at the core structure level. These complementary metrics jointly reflect the model’s ability to capture both local functional groups and global structural patterns, thereby characterizing the complexity of the generated molecular structures.
It is important to note that for all distribution-based metrics, the test set serves as the reference, while the novelty metric uses the training set as the reference to measure whether generated molecules are unseen. This distinction ensures a clear separation between distributional fidelity and generative innovation.
@@ C.2 Metrics for Molecular Structure Elucidation
The Top-$K$ Accuracy (Acc@$K$) quantifies the model’s ability to recover the exact target structure among its top-$K$ generated candidates. For each condition characterized by a set of spectra, the model is sampled $K$ times to generate $K$ molecular structure candidates. If any of the $K$ candidates exactly matches the ground-truth structure, the prediction is considered correct. Formally, the metric is defined as:
@@
Maximum Common Edge Subgraph
To quantify graph-structural similarity, we adopt the Maximum Common Edge Subgraph (MCES) distance metric:
$\operatorname{MCES}=\mathbb{E}_{\mathcal{G}}\left[|\operatorname{E}(\mathcal{G})|+|\operatorname{E}(\hat{\mathcal{G}})|-2\cdot\max_{\mathcal{H}\subseteq\mathcal{G},\,\mathcal{H}\subseteq\hat{\mathcal{G}}}|\operatorname{E}(\mathcal{H})|\right],$ (44)
where $\mathcal{G}$ and $\hat{\mathcal{G}}$ are the target and predicted molecular graphs respectively, $\operatorname{E}(\mathcal{H})$ denotes the set of edges in the common subgraph $\mathcal{H}$, and $|\operatorname{E}(\mathcal{G})|$ represents the total number of edges in graph $\mathcal{G}$. This distance metric captures the structural dissimilarity between molecular graphs, with lower values indicating greater structural similarity and better recovery of the target molecular structure.
Fingerprint-based Similarity Evaluation
@@
Fragment-based Similarity Evaluation
The Fragment-based Similarity (Fraggle) metric decomposes $\mathcal{G}$ into a set of chemically meaningful fragments $\mathcal{F}=\{f_{1},f_{2},\dots,f_{n}\}$ by strategies including double acyclic or exocyclic bond cuts with specified fragment size constraints. For each fragment $f_{i}$, Fraggle computes two Tanimoto similarities over RDKit fingerprints between the target molecule $\mathcal{G}$ and the predicted molecule $\hat{\mathcal{G}}$: one based on the standard RDKit fingerprint over the entire molecule, and another based on a masked RDKit fingerprint where atoms outside $f_{i}$ with a Tversky similarity below 0.8 are replaced with wildcards. The fragment-level score is then taken as the maximum of these two values:
@@
Functional Group-based Similarity Evaluation
[…]
$\operatorname{FGSim}=\mathbb{E}_{\mathcal{G}}\left[\frac{\left|\operatorname{FG}(\mathcal{G})\cap\operatorname{FG}(\hat{\mathcal{G}})\right|}{\left|\operatorname{FG}(\mathcal{G})\cup\operatorname{FG}(\hat{\mathcal{G}})\right|}\right$]. (50)
@@
3D Structural Evaluation
Lastly, to evaluate the accuracy of generated 3D structures, we adopt Root Mean Square Deviation (RMSD) to measure the geometric deviation between the generated 3D structures and ground-truth 3D structures. Prior to RMSD computation, atom-to-atom correspondence must be established between the generated and reference molecules. We employ the Hungarian algorithm to determine the optimal atom mapping by minimizing a weighted distance matrix. The mapping process ensures that each atom in the generated structure is paired with its most similar counterpart in the ground-truth structure based on geometric and chemical criteria. The RMSD, computed in ångströms (Å), is defined as:
$\text{RMSD}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}\|\mathbf{X}_{i}-\hat{\mathbf{X}}_{i}\|^{2}}$ (51)
where $\mathbf{X}_{i}$ and $\hat{\mathbf{X}}_{i}$ represent the 3D coordinates of the $i$-th mapped atom pair in the ground-truth and generated structures, respectively, and $n$ is the number of successfully mapped atoms.
Additionally, we compute the Atom Mapping Accuracy (MapAcc) as:
$\text{MapAcc}=\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}[\text{AtomType}(\mathbf{H}_{i})=\text{AtomType}(\hat{\mathbf{H}}_{i})]$ (52)
where $\mathbbm{1}[\cdot]$ is the indicator function, and $\text{AtomType}(\mathbf{H}_{i})$ and $\text{AtomType}(\hat{\mathbf{H}}_{i})$ denote the atom types of the $i$-th mapped atom pair, providing a complementary measure of structural fidelity that captures both geometric and chemical similarity.